OpenStudy (anonymous):

_____ √x+10=x-2 x+10 is all together under √ symbol

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i have x=0 and x=5 is this correct?

OpenStudy (lncognlto):

√(x+10)=x-2 therefore x + 10 = x^2 -4x +4 tf x^2 -5x -6 = 0 (x-6)(x+1)=0 tf x = 6 or -1.

OpenStudy (anonymous):

tf x^2 -5x -6 = 0 how u get this?

OpenStudy (lncognlto):

Subtracting x + 10 from both sides 0 = x^2 -4x +4 - x - 10 =x^2 -5x -6

OpenStudy (anonymous):

-6x=0 1x=0

OpenStudy (lncognlto):

Well, (x-6)(x+1)=0 so x-6 = 0 or x+1 = 0 so x = 6 or -1